# Measuring the (Un)Faithfulness of Concept-Based Explanations

### v3

@@ Abstract
Deep vision models perform input-output computations that are hard to interpret. Concept-based explanation methods (CBEMs) increase interpretability by re-expressing parts of the model with human-understandable semantic units, or concepts. Checking if the derived explanations are faithful—that is, they represent the model’s internal computation—requires a surrogate that combines concepts to compute the output. Simplifications made for interpretability inevitably reduce faithfulness, resulting in a tradeoff between the two. State-of-the-art unsupervised CBEMs (U-CBEMs) have reported increasingly interpretable concepts, while also being more faithful to the model. However, we observe that the reported improvement in faithfulness artificially results from either (1) using overly complex surrogates, which introduces an unmeasured cost to the explanation’s interpretability, or (2) relying on deletion-based approaches that, as we demonstrate, do not properly measure faithfulness. We propose Surrogate Faithfulness (SURF), which (1) replaces prior complex surrogates with a simple, linear surrogate that measures faithfulness without changing the explanation’s interpretability and (2) introduces well-motivated metrics that assess loss across all output classes, not just the predicted class. We validate SURF with a measure-over-measure study by proposing a simple sanity check—explanations with random concepts should be less faithful—which prior surrogates fail. SURF enables the first reliable faithfulness benchmark of U-CBEMs, revealing that many visually compelling U-CBEMs are not faithful. Code to be released.
@@ 1 Introduction
Deep learning models have delivered state-of-the-art (SOTA) results across diverse tasks, yet their internal computation remain difficult to interpret [Gerlings2020ReviewingTN], which is especially problematic in high-stakes domains such as healthcare and finance. This gap has given rise to explainable AI (XAI), whose goal is to produce explanations of model behavior. Two properties characterize explanation quality: interpretability—explanations that humans can understand—and faithfulness—explanations that reflect the model’s internal computation. XAI methods must balance the natural tension between interpretability and faithfulness. While interpretability is assessed with human studies, faithfulness must be evaluated by defining a surrogate that maps the explanation to the model’s outputs and measuring the loss between the surrogate and model. Since explanations are inevitably lossy, some discrepancy is expected. For certain XAI methods, defining a surrogate is challenging, so alternative measures, or proxies, attempt to evaluate faithfulness using different criteria.
@@
One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model [protopnet, senn]. Merging the explanation and the model eliminates the need for faithfulness checks; however, these methods often underperform black-box models by reducing model complexity (capacity) for explainability. In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution, commonly done by finding the output’s sensitivity to changes in input features, either through gradients or permutations [wang2024gradient, lime, shapley, Petsiuk2018rise, sobol]. For images, attributions are visualized as pixel-level heatmaps. Although such explanations are intuitive, developing faithfulness measures has been challenging, since many attribution methods do not specify a surrogate to reconstruct the model’s output from the explanation. Thus, proxies—such as checking if important pixels come from foreground regions or looking for significant model degradation after deleting important pixels—are used to study faithfulness. Despite strong performance on the proxies, clever sanity checks have revealed faithfulness problems [Adebayo2018SanityCF, Frye2020ShapleyManifold, Sixt2019WhenEL, Hase2021TheOP]. The resulting heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [Nguyen2021TheEO, colin2022what].
@@
Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [net2vec, netdiss, kimTCAV, zhou2014object]. To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss important aspects of the model’s computation, raising faithfulness concerns [Ramaswamy2022OverlookedFI]. To avoid these issues, unsupervised CBEMs (U-CBEMs) discover concepts automatically as concept activation vectors (CAVs)—directions in the model’s intermediate representation space—paired with a concept importance score that captures the concept’s relevance to the output. Since CBEMs operate on the model’s intermediate representations, defining a surrogate is more natural, paving the way for proper faithfulness evaluations.
@@ 2 Preliminaries
Let $E$ be an explanation function, and let $E(\mathbf{X})\in\mathcal{E}$ represent the explanation for an $\mathbf{X}\in\mathcal{X}$ (e.g., image). Let $s:\mathcal{E}\to\mathcal{Y}$ be a surrogate. Motivated by Ribeiro et al. [lime], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$. Thus, a faithfulness measure is defined by a choice of surrogate $s(.)$ and metric $d(.)$ as:
@@
U-CBEMs: To generate explanations, U-CBEMs find, for all outputs $i\in C$, a set of $K$ CAVs $V_{i}=\{\mathbf{v}_{i,k}\}_{k=1}^{K}$ and concept importances $A_{i}=\{\alpha_{i,k}\}_{k=1}^{K}$. Note that while $K$ can vary per output, all prior works find the same $K$ per output. U-CBEMs also define a mechanism $\mathcal{P}(\mathbf{h};V):\mathcal{H}\to\mathbb{R}^{K}$ to project embedding h to the concept space defined by CAVs in $V$. The explanation for output $i$ is:
@@
3 Related Works
Unsupervised Concept-Based Explanation Methods. Wang et al. [Wang2015UnsupervisedLO] find CAVs by K-Means clustering intermediate embeddings computed on a dataset. Each cluster is associated with a semantic concept (e.g., car tire, window, headlight) by tracing back the clustered embeddings to their corresponding image patches. Ghorbani et al. [Ghorbani2019ACE] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate. To ensure that the discovered CAVs fully represent the model’s behavior, Yeh et al. [ConceptSHAP] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [Young1985MonotonicSO] to find concept importance. ICE [Zhang2020InvertibleCE] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation. The authors show that concepts found with NMF are more interpretable than those found by K-Means or Principal Component Analysis (PCA). Building on this, Fel et al. [Fel2022CRAFTCR] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance. CDISCO [graziani2023concept] uses the singular value decomposition to find CAVs and a gradient-based method for finding concept importance. MCD [vielhaben2023multidimensional] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts. Finding issues with MCD’s interpretability, HU-MCD [hu-mcd] incorporates Segment Anything Model [SAM] to discover more interpretable concepts.
@@ 3 Related Works
Concept Bottleneck Models. CBEMs should not be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model. These methods [pmlr-cbm-koh20a, prob-cbm, posthoc-cbm, labelfree-cbm] train a model to first predict a set of interpretable concepts (i.e., a concept bottleneck); then, they use the concept representation for the task (e.g., object classification). By doing so, CBMs fully decompose the prediction into interpretable concepts. While CBMs are a promising approach, they typically fail to reach the same performance as black-box models, unless they introduce uninterpretable, “residual” concepts. As CBMs are inherently explainable, the notion of faithfulness does not extend to them. This work discusses faithfulness only for U-CBEMs.
@@ 4 Unifying U-CBEM Faithfulness Measures
Each U-CBEM mentioned in Sec. 3 introduces a different faithfulness measure, indicating a lack of consensus and preventing fair comparisons of faithfulness across U-CBEMs. We unify faithfulness measures under a common framework, shown in Fig. 1. Any evaluation measure compares the output of the original model with the output from the explanation (Metric). To obtain the output from the explanation, a Concept Projection transforms the intermediate representation to a concept representation, which is then passed through a Surrogate. Deletion-based proxies assert that removing the most important concepts will result in the greatest model degradation; more degradation is interpreted as a signal for higher faithfulness. Alternatively, surrogate-based measures avoid changing the input and introduce a surrogate to directly approximate faithfulness as in Eq. 3.
@@ 4.1 Deletion-Based Proxies
There are two notable issues with deletion-based proxies that prevent them from accurately evaluating faithfulness. First, it is unclear how to delete a concept. Concepts are usually deleted by Constant Deletion, commonly with a baseline of 0. This baseline must entirely delete the presence of the concept without affecting other concepts. However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [Frye2020ShapleyManifold]. Removing features by marginalization is more accurate but requires evaluating expensive, high-dimensional expectations.
Second, there are no guarantees that representations in the Deletion Space will stay on manifold after deletion. The feature attribution literature [Frye2020ShapleyManifold, akata2022manifold, taufiq2023manifold, Slack2019FoolingLA, yeh2022threading] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold. In the off-manifold regions of the data space, a highly non-linear model behaves unpredictably, so the model’s outputs may not be meaningfully related to input features. Thus, calculating feature importance with off-manifold inputs can lead to incorrect explanations [Frye2020ShapleyManifold]. More seriously, these explanations can be adversarially attacked; in [Slack2019FoolingLA], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs. Such issues can be mitigated by using on-manifold input perturbations, but this requires accurate modeling of the input distribution [Frye2020ShapleyManifold, taufiq2023manifold, vu2022emap, zaher2024manifold, Hase2021TheOP]. These problems are not unique to feature attribution; they extend to deletions in the concept and weight space, which make similar manipulations in the model’s intermediate representation space.
@@ 4.2.2 Prior Surrogate-Based Measures
Two prior works have introduced surrogate-based measures:
@@ 5.1 The SURF Surrogate
We define SURF’s surrogate $s$ to take the form of the linear layer (Eq. 9), where we use the concept representation and $A_{i}$’s obtained from any U-CBEM. The surrogate is:
@@
where $\mathcal{P}(.)_{k}$ denotes the $k$’th element. Following prior U-CBEMs and evaluation measures, SURF also operates only on the model’s final linear layer. Thus, its surrogate fully represents the model’s computation, while greatly reducing complexity compared to prior faithfulness surrogates.
@@ 5.2 The SURF Metrics
A faithful U-CBEM will have explanations with low $\text{SURF}_{\text{EMD}}$ and $\text{SURF}_{\text{MAE}}$ errors. For regression, we only use $\text{SURF}_{\text{MAE}}$.
@@ 6 Experiments
We first perform a measure-over-measure study with a basic sanity check: do faithfulness scores decrease as we progressively randomize the explanation? Only SURF passes this check. Then, we benchmark prior U-CBEMs across three tasks, revealing that SOTA methods are not faithful. Finally, we leverage SURF to analyze the trade-off between the number of concepts in an explanation and faithfulness.
@@ 6.1 Measure-over-Measure Comparison
We apply this sanity check on SURF, ICE-Eval, and C-SHAP Eval for an ImageNet-pretrained ResNet-50 [resnet] finetuned on the Caltech-101 dataset [caltech-101]. Along with the SURF metrics, we report ICE’s Top-1 Accuracy (Top-1) defined in Eq. 6 and ICE’s normalized L1 logit error (Norm L1) defined in Eq. 5. As a complementary measure to $\text{SURF}_{\text{EMD}}$, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs. Rank Corr penalizes fine-grained changes that result in a different prediction ordering but ignores differences in magnitudes. Finally, we measure surrogate complexity by its learnable parameters and FLOPs needed to obtain the output from the concept representation. Since C-SHAP-Eval’s two-layer perceptron was originally trained with cross entropy loss (CEL), we also train the surrogate with L1 loss, as this may align better with the SURF metrics. More details are in Appendix Sec. A.
Results are reported in Tab. 3, averaged over 10 seeds. In the Perfect setting, both ICE-Eval and SURF report perfect faithfulness, as expected. However, both C-SHAP-Eval variants are unable to achieve perfect faithfulness across any metric, showing C-SHAP-Eval’s limitations. In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; this occurs because their surrogates do not use concept importances, so using random concept importances has the same evaluation result as the Perfect setting. Contrast this with SURF, which clearly finds a deterioration in faithfulness across all metrics. Finally, in the Full Rand setting, C-SHAP-Eval unintuitively reports an improvement in faithfulness compared to the Perfect setting. ICE-Eval and SURF report similarly poor faithfulness scores for this setting, as expected.
Next, we turn our attention to the metrics. We highlight a failure case of Top-1 by looking at SURF’s results in the Rand Imp setting. The Top-1 score is strong; however, the other metrics (especially $\text{SURF}_{\text{EMD}}$) reveal that the reconstructed output has severe errors on the remainder of the output distribution. Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1) in the Perfect setting, despite C-SHAP-Eval (L1) directly minimizing the L1 error between logits. We attribute this inconsistency to Norm L1’s emphasis on only the groundtruth class; while the L1 error across all classes (as measured by $\text{SURF}_{\text{MAE}}$) has decreased, the normalized error for the groundtruth class increased, giving an incomplete picture of faithfulness.

### v4

@@ Abstract
Deep vision models perform input-output computations that are hard to interpret. Concept-based explanation methods (CBEMs) increase interpretability by re-expressing parts of the model with human-understandable semantic units, or concepts. Checking if the derived explanations are faithful—that is, they represent the model’s internal computation—requires a surrogate that combines concepts to compute the output. Simplifications made for interpretability inevitably reduce faithfulness, resulting in a tradeoff between the two. State-of-the-art unsupervised CBEMs (U-CBEMs) are seemingly more interpretable, while also being more faithful to the model. However, we observe that the reported improvement in faithfulness artificially results from either (1) using overly complex surrogates, which introduces an unmeasured cost to the explanation’s interpretability, or (2) relying on deletion-based approaches that, as we demonstrate, do not properly measure faithfulness. We propose Surrogate Faithfulness (SURF), which (1) replaces prior complex surrogates with a simple, linear surrogate that measures faithfulness without changing the explanation’s interpretability and (2) introduces well-motivated metrics that assess loss across all output classes, not just the predicted class. We validate SURF with a measure-over-measure study by proposing a simple sanity check—explanations with random concepts should be less faithful—which prior surrogates fail. SURF enables the first reliable faithfulness benchmark of U-CBEMs, revealing that many visually compelling U-CBEMs are not faithful. Code is released.
@@ 1 Introduction
Deep learning models have delivered state-of-the-art (SOTA) results across diverse tasks, yet their internal computation remain difficult to interpret [17], which is especially problematic in high-stakes domains such as healthcare and finance. This gap has given rise to explainable AI (XAI), whose goal is to produce explanations of model behavior. Two properties characterize explanation quality: interpretability—explanations that humans can understand—and faithfulness—explanations that reflect the model’s internal computation. XAI methods must balance the natural tension between interpretability and faithfulness. While interpretability is assessed with human studies, faithfulness must be evaluated by defining a surrogate that maps the explanation to the model’s outputs and measuring the loss between the surrogate and model. Since explanations are inevitably lossy, some discrepancy is expected. For certain XAI methods, defining a surrogate is challenging, so alternative measures, or proxies, attempt to evaluate faithfulness using different criteria.
@@
One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model [6, 3]. Merging the explanation and the model eliminates the need for faithfulness checks; however, these methods often underperform black-box models by reducing model complexity (capacity) for explainability. In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution, commonly done by finding the output’s sensitivity to changes in input features, either through gradients or permutations [47, 39, 33, 37, 10]. For images, attributions are visualized as pixel-level heatmaps. Although such explanations are intuitive, developing faithfulness measures has been challenging, since many attribution methods do not specify a surrogate to reconstruct the model’s output from the explanation. Thus, proxies—such as checking if important pixels come from foreground regions or looking for significant model degradation after deleting important pixels—are used to study faithfulness. Despite strong performance on the proxies, clever sanity checks have revealed faithfulness problems [1, 15, 41, 22]. The resulting heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [35, 7].
@@
Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [14, 4, 24, 55]. To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss important aspects of the model’s computation, raising faithfulness concerns [38]. To avoid these issues, unsupervised CBEMs (U-CBEMs) discover concepts automatically as concept activation vectors (CAVs)—directions in the model’s intermediate representation space—paired with a concept importance score that captures the concept’s relevance to the output. Since CBEMs operate on the model’s intermediate representations, defining a surrogate is more natural, paving the way for proper faithfulness evaluations.
@@ 2 Preliminaries
Let $E$ be an explanation function, and let $E(\mathbf{X})\in\mathcal{E}$ represent the explanation for an $\mathbf{X}\in\mathcal{X}$ (e.g., image). Let $s:\mathcal{E}\to\mathcal{Y}$ be a surrogate. Motivated by Ribeiro et al. [39], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$. Thus, a faithfulness measure is defined by a choice of surrogate $s(.)$ and metric $d(.)$ as:
@@
U-CBEMs: To generate explanations, U-CBEMs find, for all outputs $i\in C$, a set of $K$ CAVs $V_{i}=\{\mathbf{v}_{i,k}\}_{k=1}^{K}$ and concept importances $A_{i}=\{\alpha_{i,k}\}_{k=1}^{K}$. Note that while, for a given input, $K$ can vary per output class, all prior works find the same $K$ per output class. U-CBEMs also define a mechanism $\mathcal{P}(\mathbf{h};V):\mathcal{H}\to\mathbb{R}^{K}$ to project embedding h to the concept space defined by CAVs in $V$. The explanation for output $i$ is:
@@
3 Related Works on U-CBEMs
Wang et al. [46] find CAVs by K-Means clustering intermediate embeddings computed on a dataset. Each cluster is associated with a semantic concept (e.g., car tire, window, headlight) by tracing back the clustered embeddings to their corresponding image patches. Ghorbani et al. [18] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate. To ensure that the discovered CAVs fully represent the model’s behavior, Yeh et al. [48] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [50] to find concept importance. ICE [53] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation. The authors show that concepts found with NMF are more interpretable than those found by K-Means or Principal Component Analysis (PCA). Building on this, Fel et al. [13] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance. CDISCO [20] uses the singular value decomposition to find CAVs and a gradient-based method for finding concept importance. MCD [44] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts. Finding issues with MCD’s interpretability, HU-MCD [21] incorporates Segment Anything Model [27] to discover more interpretable concepts. Recently, sparse autoencoders (SAEs) have emerged as a promising, scalable U-CBEM for large language models, and recent attempts have applied them to interpret vision models [19, 31, 12].
CBEMs and U-CBEMs should not be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model [28, 25, 51, 36].
@@ 4 Unifying U-CBEM Faithfulness Measures
Each U-CBEM mentioned in Sec. 3 introduces a different faithfulness measure, indicating a lack of consensus and preventing fair comparisons of faithfulness across U-CBEMs. We unify faithfulness measures under a common framework, shown in Fig. 1. Any evaluation measure compares the output of the original model with the output from the explanation (Metric). To obtain the output from the explanation, a Concept Projection transforms the intermediate representation to a concept representation, which is then passed through a Surrogate. Deletion-based proxies assert that removing the most important concepts will result in the greatest model degradation; more degradation is interpreted as a signal for higher faithfulness. Alternatively, surrogate-based measures avoid changing the input and introduce a surrogate to directly approximate faithfulness as in Eq. 3. Appendix Sec. A offers a discussion and comparison with faithfulness metrics introduced for SAEs.
@@ 4.1 Deletion-Based Proxies
There are two notable issues with deletion-based proxies that prevent them from accurately evaluating faithfulness. First, it is unclear how to delete a concept. Concepts are usually deleted by Constant Deletion, commonly with a baseline of 0. This baseline must entirely delete the presence of the concept without affecting other concepts. However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [15]. Removing features by marginalization is more accurate but requires evaluating expensive, high-dimensional expectations.
Second, there are no guarantees that representations in the Deletion Space will stay on manifold after deletion. The feature attribution literature [15, 2, 43, 42, 49] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold. In the off-manifold regions of the data space, a highly non-linear model behaves unpredictably, so the model’s outputs may not be meaningfully related to input features. Thus, calculating feature importance with off-manifold inputs can lead to incorrect explanations [15]. More seriously, these explanations can be adversarially attacked; in [42], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs. Such issues can be mitigated by using on-manifold input perturbations, but this requires accurate modeling of the input distribution [15, 43, 45, 52, 22]. These problems are not unique to feature attribution; they extend to deletions in the concept and weight space, which make similar manipulations in the model’s intermediate representation space.
@@ 4.2.2 Prior Surrogate-Based Measures
Two prior works have proposed surrogate-based measures:
@@ 5.1 The SURF Surrogate
We define SURF’s surrogate $s$ to take the form of (Eq. 9), replacing $\mathbf{F}$ with the concept representation and $A_{i}$’s obtained from any U-CBEM. The surrogate is:
@@
where $\mathcal{P}(.)_{k}$ denotes the $k$’th element. Following prior U-CBEMs and evaluation measures, SURF also operates only on the model’s final linear layer. Thus, its surrogate fully represents the model’s computation, while not introducing any trainable parameters and greatly reducing complexity compared to prior faithfulness surrogates.
@@ 5.2 The SURF Metrics
We choose these specific metrics as they are easily interpretable. A faithful U-CBEM will have explanations with low $\text{SURF}_{\text{EMD}}$ and $\text{SURF}_{\text{MAE}}$ errors. For regression, we only use $\text{SURF}_{\text{MAE}}$.
@@ 6 Experiments
We first perform a measure-over-measure study with a basic sanity check: do faithfulness scores decrease as we progressively randomize the explanation? Only SURF passes this check. Then, we benchmark prior U-CBEMs across three tasks, revealing that SOTA methods are not faithful. Finally, we leverage SURF to analyze the tradeoff between the number of concepts in an explanation and faithfulness.
@@ 6.1 Measure-over-Measure Comparison
We apply this sanity check on SURF, ICE-Eval, and C-SHAP Eval for an ImageNet-pretrained ResNet-50 [23] finetuned on the Caltech-101 dataset [30]. Along with the SURF metrics, we report Top-1 Accuracy (Top-1) defined in Eq. 6 and normalized L1 logit error (Norm L1) defined in Eq. 5. As a complementary metric to $\text{SURF}_{\text{EMD}}$, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs. Rank Corr ignores magnitude differences but penalizes fine-grained changes that result in a different prediction ordering. Finally, we measure surrogate complexity by its learnable parameters and FLOPs from concept representation to output. Since C-SHAP-Eval’s surrogate was originally trained with cross entropy loss (CEL), we also train it with L1 loss, as this may align better with the SURF metrics. More details are in Appendix Sec. B.
Results are reported in Tab. 3, averaged over 10 seeds. In the Perfect setting, both ICE-Eval and SURF report perfect faithfulness, as expected. However, both C-SHAP-Eval variants do not achieve perfect faithfulness across any metric, showing C-SHAP-Eval’s limitations. In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; because their surrogates do not use concept importances, using random concept importances do not change results from the Perfect setting. Contrast this with SURF, which clearly finds a deterioration in faithfulness across all metrics. Finally, in the Full Rand setting, C-SHAP-Eval unintuitively reports an improvement in faithfulness compared to the Perfect setting. ICE-Eval and SURF report similarly poor faithfulness scores for this setting, as expected.
Next, we examine the metrics. SURF’s results in the Rand Imp setting highlight a failure case for Top-1. The Top-1 score is strong; however, the other metrics (especially $\text{SURF}_{\text{EMD}}$) reveal that the surrogate output has severe errors on the remainder of the output distribution. Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1) in the Perfect setting, despite C-SHAP-Eval (L1) directly minimizing the logit L1 error. We attribute this inconsistency to Norm L1’s emphasis on only the groundtruth class; while the L1 error across all classes (as measured by $\text{SURF}_{\text{MAE}}$) has decreased, the normalized error for the groundtruth class increased, giving an incomplete picture of faithfulness.